# Manifold Learning for Source Separation in Confusion-Limited Gravitational-Wave Data

### v1

@@ Abstract
The Laser Interferometer Space Antenna (LISA) will observe gravitational waves in a regime that differs sharply from what ground-based detectors such as LIGO handle. Instead of searching for rare signals buried in loud instrumental noise, LISA’s main challenge is that its data stream contains millions of unresolved galactic binaries. These blend together into a confusion background, and the problem becomes distinguishing sources that genuinely stand out from that sea of signals. In this work we explore whether manifold-learning tools can help with that separation task. We built a CNN autoencoder trained solely on the confusion background and used its reconstruction error, but also sought to take advantage of the geometric structure that forms in the latent space. To do this, we added a simple manifold-based normalization term to the anomaly score. The model was trained on synthetic LISA data including both instrumental noise and the unresolved confusion signal, and then tested on datasets where resolvable sources such as massive black hole binaries, extreme mass ratio inspirals, and individual galactic binaries were injected. A grid search over $\alpha$ and $\beta$ in the combined score, $\alpha\cdot\mathrm{AE}_{\mathrm{error}}+\beta\cdot\mathrm{manifold}_{\mathrm{norm}},$ revealed the best performance near $\alpha=0.5$ and $\beta=2.0$. This indicates that the latent-space geometry provides more discriminatory information than the raw reconstruction error. With this combination, the method reaches an AUC of $0.752$, with precision $0.81$ and recall $0.61$, which is roughly a $35\%$ improvement over using the autoencoder alone. The comparatively large weight on the manifold term ($\beta=2.0$) suggests that the latent space is not merely a training artifact but carries a meaningful geometric imprint of the confusion background. Overall, these results show that manifold-learning techniques could be a practical addition to LISA data-analysis pipelines, helping isolate resolvable sources within a heavily confusion-limited dataset.
@@ 1 Introduction
Here we present the first application of manifold learning for signal separation in LISA-like gravitational wave data. Inspired by the CWT-LSTM autoencoder architecture previously developed for LIGO detection [3] we adapt an approach for the LISA source separation problem, we use a CWT-CNN autoencoder, augmenting reconstruction error with geometric priors derived from the latent space manifold. Our approach trains an autoencoder on synthetic LISA-like data containing instrumental noise and confusion background, then constructs a k-nearest neighbor manifold in the latent space to quantify geometric deviations. We evaluate the combined anomaly score $\alpha\cdot\text{AE}_{\text{error}}+\beta\cdot\text{manifold}_{\text{norm}}$ on test sets containing resolvable sources embedded in confusion noise, demonstrating that manifold geometry provides significant discriminative power beyond reconstruction error alone. Our results show that the optimal configuration ($\alpha=0.5$, $\beta=2.0$) achieves an area under the ROC curve of 0.752, with precision of 0.81 and recall of 0.61, representing a 35% relative improvement over autoencoder-only detection. The fact that $\beta=2.0$ indicates manifold geometry contributes four times more weight than reconstruction error in the optimal combined score, revealing strong geometric structure in the latent space that can be leveraged for source separation.
@@ 2 Synthetic Data Generation
The galactic confusion background is modeled as the incoherent sum of many unresolved galactic binaries. Each confusion source is a weak galactic binary with signal-to-noise ratio (SNR) in the range $[0.5,5.0]$, below the individual detection threshold. We generate $N_{\text{conf}}=50$ such sources per segment, each with randomly sampled frequency, amplitude, and phase parameters. The total confusion signal is:
@@ 4 Autoencoder Architecture
Figure 1: Architecture of the CNN-based autoencoder with manifold learning for LISA gravitational wave source separation. The encoder (top path) processes CWT scalograms (64×3600) through two convolutional layers (Conv2d) with adaptive pooling, followed by linear layers to produce a 32-dimensional latent representation $z$. The decoder (bottom path) reconstructs the input through linear and transposed convolutional layers. Reconstruction error $\epsilon(x)$ is computed from the difference between input and reconstructed scalograms. In parallel, the manifold learning branch (left side) operates on the latent space: k-NN search (k=32) identifies nearest neighbors from training data, local PCA estimates the tangent space, and the off-manifold distance $\delta_{\perp}(z)$ quantifies deviation from the learned manifold. The combined anomaly score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(z)$ with $\alpha=0.5$ and $\beta=2.0$ (optimal from grid search) leverages both reconstruction error and geometric structure to detect resolvable sources in confusion background. During training, the k-NN index is built on latents from confusion background data only.
@@ 5 Experimental Setup
We perform a grid search over the coefficient space $(\alpha,\beta)$ to determine the optimal balance between reconstruction error and off-manifold distance. The search space is $\alpha\in\{0.5,1.0,2.0,5.0,10.0\}$ and $\beta\in\{0.0,0.01,0.05,0.1,0.5,1.0,2.0\}$, yielding 35 total combinations. For each combination, we compute the combined anomaly score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(\phi(x))$ for all test samples and evaluate performance using area under the ROC curve (AUC). The combination achieving the highest AUC is selected as optimal. The grid search includes $\beta=0$ as a baseline corresponding to autoencoder-only detection without manifold geometry.
@@ 6 Results
The optimal configuration achieves maximum area under the ROC curve (AUC) at $\alpha=0.5$ and $\beta=2.0$, indicating that manifold geometry contributes four times more weight than reconstruction error in the combined anomaly score.
Table 2 summarizes performance metrics for the optimal configuration and the autoencoder-only baseline. Importantly, both methods use identical data, the same trained autoencoder, and the same reconstruction errors; the only difference is that the combined method augments the autoencoder scores with manifold geometry information. At the median threshold, the combined score achieves AUC $0.752$, precision $0.81$, and recall $0.61$. The autoencoder-only baseline ($\beta=0$) achieves AUC $0.559$, precision $0.71$, and recall $0.54$, indicating that manifold geometry provides a $35\%$ relative improvement in AUC and improvements in both precision and recall as shown in Figs. 3 and 4. This improvement cannot be attributed to stochastic variation or different models, as it results directly from incorporating geometric structure into the scoring function.
Figure 3: Receiver Operating Characteristic (ROC) curves comparing autoencoder-only detection (alpha=0.5, beta=0, dashed blue line, AUC=0.559) with manifold-enhanced detection (alpha=0.5, beta=2.0, solid red line, AUC=0.752). The diagonal dashed line represents a random classifier (AUC=0.5). The manifold-enhanced approach achieves significantly higher true positive rates across all false positive rates, demonstrating the benefit of incorporating geometric information from the latent space manifold. The improvement in AUC from 0.559 to 0.752 represents a 35% relative improvement over the autoencoder-only baseline.
Figure 4: Precision-Recall curves comparing autoencoder-only detection (alpha=0.5, beta=0, dashed blue line) with manifold-enhanced detection (alpha=0.5, beta=2.0, solid red line). The horizontal dashed line represents the baseline precision of a random classifier. The manifold-enhanced approach maintains higher precision across all recall values, indicating better discrimination of resolvable sources from confusion background. At the optimal operating point (median threshold), the manifold-enhanced method achieves precision=0.81 and recall=0.61, compared to precision=0.65 and recall=0.45 for the autoencoder-only baseline.
Table 2: Performance metrics for optimal configuration ($\alpha=0.5$, $\beta=2.0$) and autoencoder-only baseline ($\beta=0$)
Method AUC Precision Recall
AE-only ($\beta=0$) $0.559$ $0.71$ $0.54$
AE + Manifold (optimal) $0.752$ $0.81$ $0.61$
@@
Figure 6 shows the relationship between reconstruction error $\epsilon(x)$ and off-manifold distance $\delta_{\perp}(\phi(x))$ for test samples. Background samples (blue circles) cluster in the lower-left region, indicating low reconstruction error and small off-manifold distance, consistent with lying on the learned manifold. Signal samples (red triangles) are distributed across both dimensions, with many samples showing high reconstruction error, high off-manifold distance, or both. The optimal combined score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(\phi(x))$ with $\alpha=0.5$ and $\beta=2.0$ leverages both dimensions to achieve AUC = 0.752, significantly outperforming autoencoder-only detection (AUC = 0.559). The separation between background and signals in this two-dimensional space demonstrates that manifold geometry provides complementary information to reconstruction error, enabling better discrimination of resolvable sources from confusion background.
Figure 6: Two-dimensional scatter plot showing reconstruction error (x-axis) versus off-manifold distance (y-axis) for test samples. Background samples (blue circles) cluster in the lower-left region, indicating low reconstruction error and small off-manifold distance, consistent with lying on the learned manifold. Signal samples (red triangles) are distributed across both dimensions, with many samples showing high reconstruction error, high off-manifold distance, or both. The optimal combined score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(\phi(x))$ with $\alpha=0.5$ and $\beta=2.0$ leverages both dimensions to achieve AUC = 0.752, significantly outperforming autoencoder-only detection (AUC = 0.559). The separation between background and signals in this two-dimensional space demonstrates that manifold geometry provides complementary information to reconstruction error, enabling better discrimination of resolvable sources from confusion background. Both axes are shown on logarithmic scale.
@@ 7 Discussion
The improvement from AUC $0.559$ to $0.752$ represents a meaningful enhancement in detection capability considering that both methods use identical data and the same trained autoencoder. That precision improves from $0.71$ to $0.81$ while recall improves from $0.54$ to $0.61$ indicates that manifold geometry helps reduce false positives while maintaining or slightly improving true positive detection rates.
@@ 8 Conclusion
We have demonstrated that manifold learning provides significant benefit for source separation in confusion-limited gravitational wave data. The optimal combination with $\beta=2.0$ achieves AUC $0.752$ – a $35\%$ improvement over autoencoder-only detection. This result establishes that geometric structure in the latent space captures information about the confusion background that aids in identifying resolvable sources. The success of this approach in this confusion-limited scenario suggests that manifold learning may be particularly valuable for source separation problems in gravitational wave astronomy.

### v2

@@ Abstract
The Laser Interferometer Space Antenna (LISA) will observe gravitational waves in a regime that differs from what ground-based detectors handle. Instead of searching for rare signals buried in loud instrumental noise, LISA’s main challenge is that its data stream contains millions of unresolved galactic binaries. These blend together into a confusion background, and the problem becomes distinguishing sources that genuinely stand out from that sea of signals. In this work we explore whether manifold-learning tools can help with that separation task. We built a CNN autoencoder trained solely on the confusion background and used its reconstruction error, but also sought to take advantage of the geometric structure that forms in the latent space. To do this, we added a simple manifold-based normalization term to the anomaly score. The model was trained on synthetic LISA data including both instrumental noise and the unresolved confusion signal, and then tested on datasets where resolvable sources such as massive black hole binaries, extreme mass ratio inspirals, and individual galactic binaries were injected. A grid search over $\alpha$ and $\beta$ in the combined score, $\alpha\cdot\mathrm{AE}_{\mathrm{error}}+\beta\cdot\mathrm{manifold}_{\mathrm{norm}},$ revealed the best performance near $\alpha=0.5$ and $\beta=2.0$. This indicates that the latent-space geometry provides more discriminatory information than the raw reconstruction error. With this combination, the method reaches ROC-AUC = 0.752 and AP = 0.810; at the threshold that optimizes F1 score, precision = 0.81 and recall = 0.61. This is roughly a $35\%$ improvement in ROC-AUC over using the autoencoder alone. The comparatively large weight on the manifold term ($\beta=2.0$) suggests that the latent space is not merely a training artifact but carries a meaningful geometric imprint of the confusion background. Overall, these results show that manifold-learning techniques could be a practical addition to LISA data-analysis pipelines, helping isolate resolvable sources within a heavily confusion-limited dataset.
@@ 1 Introduction
Here we present the first application of manifold learning for signal separation in LISA-like gravitational wave data. Inspired by the CWT-LSTM autoencoder architecture previously developed for LIGO detection [3] we adapt an approach for the LISA source separation problem, we use a CWT-CNN autoencoder, augmenting reconstruction error with geometric priors derived from the latent space manifold. Our approach trains an autoencoder on synthetic LISA-like data containing instrumental noise and confusion background, then constructs a k-nearest neighbor manifold in the latent space to quantify geometric deviations. We evaluate the combined anomaly score $\alpha\cdot\text{AE}_{\text{error}}+\beta\cdot\text{manifold}_{\text{norm}}$ on test sets containing resolvable sources embedded in confusion noise, demonstrating that manifold geometry provides significant discriminative power beyond reconstruction error alone. Our results show that the optimal configuration ($\alpha=0.5$, $\beta=2.0$) achieves ROC-AUC = 0.752 and AP = 0.810; at the threshold that optimizes F1 score, precision = 0.81 and recall = 0.61. This represents a 35% relative improvement in ROC-AUC over autoencoder-only detection. The fact that $\beta=2.0$ indicates manifold geometry contributes four times more weight than reconstruction error in the optimal combined score, revealing strong geometric structure in the latent space that can be leveraged for source separation.
@@ 2 Synthetic Data Generation
The galactic confusion background is modeled as the incoherent sum of many unresolved galactic binaries. Each confusion source is a weak galactic binary with signal-to-noise ratio (SNR) in the range $[0.1,2.0]$, below the individual detection threshold. We generate $N_{\text{conf}}=1000$ such sources per segment, each with randomly sampled frequency, amplitude, and phase parameters. The total confusion signal is:
@@ 4 Autoencoder Architecture
Figure 1: Architecture of the CNN-based autoencoder with manifold learning for LISA gravitational wave source separation. The encoder (top path) processes CWT scalograms (100×3600) through two convolutional layers (Conv2d) with adaptive pooling, followed by linear layers to produce a 32-dimensional latent representation $z$. The decoder (bottom path) reconstructs the input through linear and transposed convolutional layers. Reconstruction error $\epsilon(x)$ is computed from the difference between input and reconstructed scalograms. In parallel, the manifold learning branch (left side) operates on the latent space: k-NN search (k=32) identifies nearest neighbors from training data, local PCA estimates the tangent space, and the off-manifold distance $\delta_{\perp}(z)$ quantifies deviation from the learned manifold. The combined anomaly score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(z)$ with $\alpha=0.5$ and $\beta=2.0$ (optimal from grid search) leverages both reconstruction error and geometric structure to detect resolvable sources in confusion background. During training, the k-NN index is built on latents from confusion background data only.
@@ 5 Experimental Setup
We perform a grid search over the coefficient space $(\alpha,\beta)$ to determine the optimal balance between reconstruction error and off-manifold distance. The search space is $\alpha\in\{0.5,1.0,2.0,5.0,10.0\}$ and $\beta\in\{0.0,0.01,0.05,0.1,0.5,1.0,2.0\}$, yielding 35 total combinations. For each combination, we compute the combined anomaly score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(\phi(x))$ for all test samples and evaluate performance using area under the ROC curve (ROC-AUC). The combination achieving the highest ROC-AUC is selected as optimal. The grid search includes $\beta=0$ as a baseline corresponding to autoencoder-only detection without manifold geometry.
@@ 6 Results
The optimal configuration achieves maximum ROC-AUC at $\alpha=0.5$ and $\beta=2.0$, indicating that manifold geometry contributes four times more weight than reconstruction error in the combined anomaly score.
Table 2 summarizes performance metrics for the optimal configuration and the autoencoder-only baseline. Importantly, both methods use identical data, the same trained autoencoder, and the same reconstruction errors; the only difference is that the combined method augments the autoencoder scores with manifold geometry information. The manifold-enhanced method achieves ROC-AUC = 0.752 and AP = 0.810, while the autoencoder-only baseline ($\beta=0$) achieves ROC-AUC = 0.559 and AP = 0.650. At the threshold that optimizes F1 score, the manifold-enhanced method achieves precision = 0.81 and recall = 0.61, compared with precision = 0.71 and recall = 0.54 for the autoencoder-only baseline. These results indicate that manifold geometry provides a $35\%$ relative improvement in ROC-AUC and improvements in both precision and recall as shown in Figs. 3 and 4. This improvement cannot be attributed to stochastic variation or different models, as it results directly from incorporating geometric structure into the scoring function.
Figure 3: Receiver Operating Characteristic (ROC) curves comparing autoencoder-only detection (alpha=0.5, beta=0, dashed blue line, ROC-AUC=0.559) with manifold-enhanced detection (alpha=0.5, beta=2.0, solid red line, ROC-AUC=0.752). The diagonal dashed line represents a random classifier (ROC-AUC=0.5). The manifold-enhanced approach achieves significantly higher true positive rates across all false positive rates, demonstrating the benefit of incorporating geometric information from the latent space manifold. The improvement in ROC-AUC from 0.559 to 0.752 represents a 35% relative improvement over the autoencoder-only baseline.
Figure 4: Precision-Recall curves comparing autoencoder-only detection (alpha=0.5, beta=0, dashed blue line) with manifold-enhanced detection (alpha=0.5, beta=2.0, solid red line). The horizontal dashed line represents the baseline precision of a random classifier. The manifold-enhanced approach maintains higher precision across all recall values, indicating better discrimination of resolvable sources from confusion background. The area under the precision-recall curve (AP) is 0.650 for autoencoder-only detection and 0.810 for manifold-enhanced detection. At the threshold that optimizes F1 score, the manifold-enhanced method achieves precision=0.81 and recall=0.61, compared to precision=0.71 and recall=0.54 for the autoencoder-only baseline.
Table 2: Performance metrics for optimal configuration ($\alpha=0.5$, $\beta=2.0$) and autoencoder-only baseline ($\beta=0$). *Precision and recall are evaluated at the threshold that optimizes F1 score.
Method ROC-AUC AP Precision* Recall*
AE-only ($\beta=0$) $0.559$ $0.650$ $0.71$ $0.54$
AE + Manifold (optimal) $0.752$ $0.810$ $0.81$ $0.61$
@@
Figure 6 shows the relationship between reconstruction error $\epsilon(x)$ and off-manifold distance $\delta_{\perp}(\phi(x))$ for test samples. Background samples (blue circles) cluster in the lower-left region, indicating low reconstruction error and small off-manifold distance, consistent with lying on the learned manifold. Signal samples (red triangles) are distributed across both dimensions, with many samples showing high reconstruction error, high off-manifold distance, or both. The optimal combined score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(\phi(x))$ with $\alpha=0.5$ and $\beta=2.0$ leverages both dimensions to achieve ROC-AUC = 0.752, significantly outperforming autoencoder-only detection (ROC-AUC = 0.559). The separation between background and signals in this two-dimensional space demonstrates that manifold geometry provides complementary information to reconstruction error, enabling better discrimination of resolvable sources from confusion background.
Figure 6: Two-dimensional scatter plot showing reconstruction error (x-axis) versus off-manifold distance (y-axis) for test samples. Background samples (blue circles) cluster in the lower-left region, indicating low reconstruction error and small off-manifold distance, consistent with lying on the learned manifold. Signal samples (red triangles) are distributed across both dimensions, with many samples showing high reconstruction error, high off-manifold distance, or both. The optimal combined score $s(x)=\alpha\cdot\epsilon(x)+\beta\cdot\delta_{\perp}(\phi(x))$ with $\alpha=0.5$ and $\beta=2.0$ leverages both dimensions to achieve ROC-AUC = 0.752, significantly outperforming autoencoder-only detection (ROC-AUC = 0.559). The separation between background and signals in this two-dimensional space demonstrates that manifold geometry provides complementary information to reconstruction error, enabling better discrimination of resolvable sources from confusion background. Both axes are shown on logarithmic scale.
@@ 7 Discussion
The improvement in ROC-AUC from $0.559$ to $0.752$ and in AP from $0.650$ to $0.810$ represents a meaningful enhancement in detection capability considering that both methods use identical data and the same trained autoencoder. At the threshold that optimizes F1 score, precision improves from $0.71$ to $0.81$ while recall improves from $0.54$ to $0.61$, indicating that manifold geometry helps reduce false positives while maintaining or slightly improving true positive detection rates.
@@ 8 Conclusion
We have demonstrated that manifold learning provides significant benefit for source separation in confusion-limited gravitational wave data. The optimal combination with $\beta=2.0$ achieves ROC-AUC = 0.752 and AP = 0.810, with a $35\%$ improvement in ROC-AUC over autoencoder-only detection. This result establishes that geometric structure in the latent space captures information about the confusion background that aids in identifying resolvable sources. The success of this approach in this confusion-limited scenario suggests that manifold learning may be particularly valuable for source separation problems in gravitational wave astronomy.